Thornton Industries began construction of a warehouse on July 1, 2018. The project was completed on March 31, 2019. No new loans were required to fund construction. Thornton does have the following two interest-bearing liabilities that were outstanding throughout the construction period: $3,000,000, 6% note $5,000,000, 2% bonds Construction expenditures incurred were as follows: July 1, 2018 $ 540,000 September 30, 2018 750,000 November 30, 2018 750,000 January 30, 2019 690,000 The company’s fiscal year-end is December 31. Required: Calculate the amount of interest capitalized for 2018 and 2019.
Answer #1

Interest to be capitalize the lesser of Actual interest cost and Avoidable interest.

1. Amount of interest capitalized for 2018:

i. Calculation of Weighted Average Accumulated expenditures:

Date Actual
Expenditure
x Capitalization
Period
= Weighted Average
Accumulated expenditures
07/01/2018 $540,000 x 6/12 = $270,000
09/30/2018 $750,000 x 3/12 = $187,500
11/30/2018 $750,000 x 1/12 = $62,500
$2,040,000 $520,000

ii.Calculation of Actual interest:

Debt x Interest
Rate*
= Actual
Interest
$3,000,000 x 3% = $90,000
$5,000,000 x 1% = $50,000
$8,000,000 $140,000

*Interest rate for 6 months

Weighted-Average interest rate on general debt = 140,000/8,000,000 = 1.75%

iii. Calculation of Avoidable interest:

Accumulated
Expenditures
x Interest
Rate*
Avoidable
Interest
520,000 x 1.75% $9,100

*Interest rate is Weighted-Average interest rate which is computed above.

Amount to be capitalized: $9,100 because it is lesser.

.

2. Amount of interest capitalized for 2019:

i. Calculation of Weighted Average Accumulated expenditures:

Date Actual
Expenditure
x Capitalization
Period
= Weighted Average
Accumulated expenditures
01/30/2018 $690,000 x 2/3 = $460,000

ii.Calculation of Actual interest:

Debt x Interest
Rate
= Actual
Interest
$3,000,000 x 1.50% = $45,000
$5,000,000 x 0.50% = $25,000
$8,000,000 $70,000

*Interest rate for 3 months

Weighted-Average interest rate on general debt = 70,000/8,000,000 = 0.875%

iii. Calculation of Avoidable interest:

Accumulated
Expenditures
x Interest
Rate
Avoidable
Interest
$460,000 x 0.875% $4,025

*Interest rate is Weighted-Average interest rate which is computed above.

Amount to be capitalized: $4,025 because it is lesser.